CH/Meet/93/02

Minutes

Charissa Group Meeting

Meeting held on: 7 April 1993

Meeting held at: Daresbury

Present:

BRF(Chair),WNC,GJG,SJAH,RAH,JK, JSL,ACM,VP,WDMR,RPW,DLW

Apologies:  NMC,RAC,CDJ,GT,JZ
Distribution: Those above plus DDW, all by e-mail/file-server.

Minutes

1. Minutes of Meeting of 10 February 1993: accepted

2. Matters Arising:

JK requested that he and D.Warner be added to the distribution list of the Charissa minutes. JK stayed at the present meeting for Items 1,2,5,6 and 9.

BRF reported that Charissa labels for electronics modules would be arranged

and afixed upon unpacking in Australia.  SJAH

3. Update on NSC/SERC matters:

Present funding situation: there was still no decision on the ABRC flexibility funds required to fund nuclear physics from 1996/97. The government’s White Paper was awaited.

RAL Test Bed: this funding had not yet been authorised by SERC. [N.B. This authorisation has occurred since the Meeting.]

Following the closure of the NSF: the disposal of minor equipment from DL would be reviewed again by DL staff and spokespersons for the NSC programmes. All equipment would be issued on loan and could be requested by other groups.

The next meeting of the joint French/UK management committee dealing with Eurogam would also include Charged Particle representatives.

The support group at DL was likely to be very heavily loaded with design and drafting requests, and effort must be planned well in advance. This was discussed, and it was agreed that outline plans for the new scattering chamber

were required by September 1993. (See also, Item 5.)  JSL,DLW,JK

The support group staff will be funded through DL for travel in support of core activities, but meetings to discuss possible future developments may require funding from University rolling grant equipment lines.

Future SERC committee structure: the NSC will be superceded by the Nuclear Physics Committee (NPC) of the new board for Particles, Astronomy and Space.

A Charged Particle User Group is formally required by the NPC. It was agreed by the meeting that its present membership provided suitable representation of these interests.

Allocation of Beamtime at the ANU:

A joint NPC/ANU panel will approve programmes proposed for the ANU. Proposals will be submitted at any time, and the physics case and feasibility will be considered. Recommendations will be passed to the NPC, whose approval will be valid for 18 months. For running in April and September 1994, this implies submissions in August 93 and February 1994 in practice. Scheduling of beamtime will be done by the ANU, who require at least 12 weeks’ notice.

A Charissa programme proposal is required before August.  BRF

Funding of Travel to ANU Experiments:

Travel costs for Staff will need to be justified and supported, and charged against the grant. It was planned to continue supporting travel to the ANU via the university grants.

Students will be funded through the Studentship office, who will decide what fraction of the requested costs they will award. Bids are made in advance and payment is retrospective (to a UK bank account) on production of supporting documentation and receipts. Each group should note that SERC-provided insurance for students will cover medical and baggage insurance, but death and disability insurance must be provided by other means.

Regarding student funding, the meeting was concerned at the variations between the treatment of similar claims. It was agreed that BRF should prepare wording for bids to cover AM and ML, who are expected to be travelling out

first.  BRF

Annual Review of Programs:

It was noted that the NPC will require a review of the current Charissa programme at its meeting in September 1993, where it will grade the various programmes. It was expected that this would comprise a 30 minute presentation, supported by a short written summary and statistical details submitted in advance. It was thought that the review would concentrate on scientific development and would be required to highlight successes and problems. It should also address the continued participation of groups, the financial statements and planned expenditure, use of DL minor and major equipment, requests for same by other groups, administrative concerns, and a statement of DL and university manpower.

It was agreed that each staff member should provide input to BRF towards

the preparation of the review.  WNC,NMC,JSL,WDMR,DLW

It was further agreed that a meeting should be held to help prepare for the

review, with a representative from each university.  BRF

4. Progress on Detector Tests:

DLW reported results obtained by Steve Chappell from beam tests of detectors under assessment:

(a) Sheet resistance 2–D PSD’s Two detectors were tested with position signals taken from each corner through 510Ω resistors, and an energy signal taken from the back. The beam was carbon on a gold target, with detectors near 20 degrees (lab). A new calibration grid was employed, with wires of 0.5, 0.3, 0.2 and 0.1 mm diameter.

i. Intertechnique supplied a detector as used by the MULTICS collaboration. This displayed the familiar pin cushion effect. A simple 1–D projection of uncorrected data showed the 0.2 mm wires.

ii. Hamamatsu supplied (for £3K) a detector with a lower-resistance on the perimeter of the resistive surface. A pepperpot mask gave an undistorted spectrum on-line. The wire grid gave uncorrected 1–D projections in which the 0.1 mm wires were very clearly seen. The energy signal depends strongly on position for short time constants (20% dip at centre for 1μs, ≈0% dip for 10μs). There was more data to be analysed for intermediate time constants. A single hole in the pepperpot gave 230 keV resolution for 85 MeV 12C ions scattered from 100 μg cm^{−2} gold.

(b) Strip Detectors Detectors supplied by Hamamatsu appeared to give energy and position resolution as good or better than Micron detectors, at 30% less cost. The Hamamatsu detectors also appeared to be much less prone to deterioration in-beam. The leakage was constant at ∼ 1μA while a Micron detector suffered an increase from 1 to 40 μA.

A full report and decision regarding purchase of the 2–D detectors for the new array should be completed by September.

5. Progress on Chamber and Detector Assemblies:

JK reported that prototypes were under construction at York. A vessel for nine telescopes was completed, and the backplate nearly completed. After discussions with NMC, the retaining clamp for the scintillators had been altered to be Tufnol and the springs would have extra tension. The Verelec vacuum feedthroughs were due at the end of April. Preamp can drawings had been issued to York and Oxford for information on 6 April. GA’s for the nine-telescope module and beam-line layout would still require significant work.

Anode and electrode drawings were with RAC for final comment.  RAC
NMC had requested a second scintillator mount.  DLW

It was agreed that JK should order parts required for prototype construction

through York.  JK,DLW

Timescales: it was agreed that a module should be ready for tests at the ANU in March 1994. Availability for testing in September 1993 was attractive, but not essential. A full assembly should be ready in time for running in September 1994. These timescales were considered to be realistic.

Priorities: The following were agreed during the meeting, under the indicated

Agenda items:  JK

Table 1: Priorities for Workshop Jobs

Date Required Description of Job  Agenda
Immediate Modified detector arm plates 9
Flange adaptor  10

Other jobs for next run (Sep 93)

August 93 Possible mods to arm plates 9
September 93 Outline plans for new chamber  3
Mar 94 at ANU Nine-telescope module  5
Sep 94 at ANU New gas handling system  6
Sep 94 at ANU Full MEGHA assembly for running  5

6. Progress on Gas Supply:

Design details had been agreed at the most recent MEGHA meeting and construction was proceeding at York.

JK requested clarification of timescales. The new system was required for operation at the ANU in September 1994. (cf. Item 5, above.)

RAC and JK would visit York to assess progress, which was believed to be

good.  RAC,JK

7. Progress on Electronics and Data Acquisition:

Preamplifiers and Bias Supplies; Oxford Manpower:

WDMR reported that the design was essentially complete.

The Department at Oxford had now implemented fully the system where technical staff are supported by grants. Charissa has effort from RAH (×1.0) and Dennis Savory (×0.5). It is certain that the manufacture of preamplifiers and biases will need to be contracted out. WDMR also stressed that tests on the completed items, and organization of these tests, would require help from outside Oxford. York was mentioned as a possible contributor.

Modifications to Event Manager EM4:

P. Simmons discussed this, with reference to document Ch/Tech/93/06. The aim was to modify the CAMAC to VME transfer, in order to release VME processor power for sorting and speed the transfer.

In the present system, there is a serial link from the Read-and-Store module, via the Crate Controller, to the GEC emulator in VME. The serial CAMAC to VME transfer is 20 times slower than the HSM transfer rate.

In the new system, a Starburst in CAMAC would respond to a LAM raised by the Event Control Module: it would read the header and the appropriate CA-MAC parameters (TDC’s, etc.) and write them to the base of the Data Stack (ADCs already written directly to stack by FERA transfer). The existing CAMAC data stack (256 words capacity) would then output the completed event via FERA to an HSM unit in VME.

The new system involves changing the header format for events. The implications for the on-line sorting program need to be investigated and the necessary

modifications made.  VP

P. Simmons will write the necessary Starburst code in BASIC, and first tests will use this. RAH estimated that the modifications to the ECM unit would take 2 weeks.

An assembler version of the BASIC code was needed for data acquisition. (In assembler, a CAMAC read is 100 times faster.) BRF expressed concern at the drain on manpower that this task could involve. After discussion, it was agreed that a contractor familiar with the system should be asked to tender.

RAH agreed to approach Golding.  RAH

Charissa Amplifier Development:

RAH reported, with reference to CH/Tech/93/05.

Each amplifier has a Channel Control Unit that can remotely trim the fine gain, pole zero, and discriminator level. All settings are still available at front panel.

Regarding the charge-to-time conversion, three prototypes are presently being assessed at Oxford, particularly with regard to linearity and the performance of surface-mount chips. These tests will be followed by tests for gain drifts etc. with gamma sources, by P. Simmons.

Construction is virtually complete at Oxford of 32 new amplifiers that have been packaged into half-width modules. Construction of 32 prototype channel controls is also proceeding. The amplifiers that will be used in September at the ANU will be modified at ANU after the experiment.

New amplifiers (256 in addition to above) are planned using surface-mounted components. To go with these, 256 channel controls are also planned. Costings should be obtained before the July Charissa meeting. Daresbury should be asked to tender for surface-mounted manufacture if pos-

sible.  WDMR,RAH

Oxford should liaise with York electronics regarding the assembly and testing of the new amplifiers – the tests will need to be quite sophisticated. DLW

reported that one person equivalent could possibly be provided.  RAH

Oxford and York should liaise with Daresbury regarding a possible contribu-

tion by DL to the amplifier assembly.  WDMR,DLW

8. Developments on Event Manager and GEC Emulator:

VP noted that the last two scheduled NSF experiments by Charissa had used the new Emulator ^{∗} system for data acquisition. This had been successful.

There had been some inconvenience due to the transfer of data from the event manager to VME, which slowed down the on-line sorting but had been anticipated. The discussion under the previous Agenda item included the most recent work to modify the current Event Manager.

VP reported that Hytec have delivered the unit ordered by Charissa for the VME/CAMAC interface, and the EPROM to repair problems in the first one.

The coding for set up and reading of the HSM module in VME is underway at DL. A copy has been given to SJAH.

9. Requirements for Hybrid Detector Operation at the ANU:

BRF reported that the Planning Trip to ANU had identified the need for a revised method of detector mounting. It was suggested that an existing DL arm plate would be mounted upside down, with the hybrids suspended. S.Metcalf would work on this, and the scheme would be discussed at a meeting on 30 April.

The new mounting plate was required at ANU for tests in July. Any significant

modifications could be made in August, in the UK if required.  JK

This modified mount has the highest priority (see Item 5).

The gas handling for the forthcoming September run will use the system already in place at the ANU (where propane is now used).

∗VP recommends that the phrase “GEC emulator” be replaced by “MVME187” or “data acquisition processor”

10. Report on Planning Trip to the ANU:

BRF presented a document summarising the actions arising from the planning trip to ANU by BRF and JSL, in the previous week. ^{†}

Unpacking of Container: we are grateful that David Weisser has agreed to oversee arrival and unpacking of the container.

Office Space: an office with space for 3–4 will be provided.

Visa Applications: BRF will circulate application forms and a letter from

ANU.  BRF

Local Equipment and Incidental Expenses: a $10K fund would be set up at ANU via Birmingham. Charges to this fund will be authorised by G. Dracoulis via Paddy Regan. This fund will allow, e.g., materials from Stores and work by School Electronics to be paid for. ($10K ≈ £5K).

Magnesium Beam Tests: a 170 MeV ^{24}Mg beam would be tested at ANU within the next 6 weeks, and we would be advised of the results.

Scattering Chamber Details: The ‘Charissa at ANU’ document Ch/Doc/93/04 contains a list of 21 points concerning chamber and cable installation, including the following.

The arm readouts were limited to 0.1 degree accuracy, and ANU are pursuing increased accuracy. The decision at the previous Charissa meeting, to buy extra carriages for inside the chamber, was revoked because their fitting would not be easy to achieve.

It is required that the shielding on the signal cables be isolated from the chamber, so the details of the feedthroughs on the Daresbury chamber should be checked. An adaptor from the existing Charissa flanges to 4-inch Dependex

(used at ANU) must be made. This is a top priority.  DLW,JK

Details of the proposed arrangement of cable racks, preamps and gas control around the chamber are included in Ch/Doc/93/04.

Targets: can be made at ANU. Standard carbon and gold thicknesses, and flash gold have been requested already.

†This handwritten document was circulated at the meeting, and has been allocated the number Ch/Doc/93/04, ‘Charissa at ANU’. More details may be found there.

11. Timetabling of Manpower for the ANU Experiment:

The proposals tabled by BRF in the previous Item involved three phases in the running of the experiment:

Installation Phase: install chamber mounts and feedthroughs, main cable run, control room racks, electronics and DA computers.

Commissioning Phase: commission electronics and DA system, including two test runs with beam.

Running Phase: two periods of running, 27 Sep – 4 Oct as two 4–day runs, 7–day break, then 11 – 26 Oct as three 4–day runs followed by clearup of chamber and electronics.

The suggested timings, after discussions with the ANU, are given in Table 2.

Table 2: Suggested Manpower for Next Run

Phase  Duration From→ Until Core  Other
Installation  6 weeks 5/7 → 13/8 SJAH  BRF,GT,AM,ML
Commissioning 4 weeks 30/8 → 24/9 SJAH,RPW,

CDJ

Running  5 weeks 27/9 →29/10 SJAH,RPW, NMC,WNC,GJG,
CDJ  JP,NC,SS

followed by

WDMR,DLW,MF, PS,SF,SC

Generally, these were endorsed by the meeting. WDMR indicated that there should be some overlap in time between the two different groups covering the running phase, in addition to the continuing core group. DLW raised the need for contingency plans, should the shipment of equipment be delayed.

Accommodation was recommended at Liversidge Apartments with overflow at University House, both located on campus and reasonably priced ($357/wk for two-bed apartment, and $42/day pp respectively, $ ≈£×2.05). Precise dates

will be discussed at the next meeting, and firm bookings decided.  ALL

12. Packing for Shipment to the ANU:

Most of the electronics and data acquisition system had been packed up on 5/6 April by SJAH, RPW and RAH.

The Suns ordered for use at ANU would be tested at DL prior to shipping in original packing. One exabyte drive presently at Birmingham and one at DL are to go. The SparcPrinter is to be tested and packed along with toner cartridges, but the interface cards will be kept at DL for testing. VP,SJAH

Packing for shipment will start on 19 April, with Union Air Transport Company. A note describing the general procedure had been circulated by DLW in the previous week (Ch/Doc/93/05). DLW would be coordinating, and it was agreed that JSL and SJAH would also be present on the 19th. DLW,JSL,SJAH

Computer manuals and exabyte tapes, hybrid detectors, cable runs and patch panels would be packed. Designated people would ensure that the target-area patch panels and long cable runs to the vault room (DLW/JK) and countingarea patch panels (JSL/WNC) were packed. JSL suggested that HV cables should be included for bias supplies.

Extra connectors and a crimping tool for the blue superscreen cable should

be ordered.  WNC,GJG

BRF planned to visit DL on 20 April to try and finalise DL minor details

relevant to the packing.  BRF
The final day for packing was reported to be Friday 30 April.  ALL

13. Documentation for ANU Running:

The document ‘Charissa at ANU’ cited above includes helpful comments on documentation (Ch/Doc/93/04).

The values of the various electronics modules being shipped to ANU were discussed for purposes of insurance. The value of the data acquisition system required clarification.

The latter information should be forwarded to SJAH.  RAH

The customs procedure – in particular, whether part of the shipment can

easily be brought back, independently – needs to be checked.  BRF

14. Plans for Strasbourg:

NMC and DLW will be giving invited talks about Charissa at the Heavy Ion Workshop being held in Strasbourg on 5–7 May.

Two representatives each from Surrey and York, one from Manchester (JSL!), and several from Birmingham will be attending. It was hoped to build on earlier contacts with the ICARE group and others at Strasbourg, with a view to future collaboration.

JSL agreed to investigate shipping the existing Charissa chamber to Stras-

bourg, as agreed at the previous Charissa meeting.  JSL

15. Any Other Business:

(a) A letter had been received by the Universities, from DL, regarding money that had been brought forward to buy Charissa equipment. This would now be charged against the Grants, as previously agreed. It was agreed that at each university, the equipment line in the grant should have the Charissa component clearly separated. WNC agreed to ensure that the funds brought forward were debited against computing

and not Charissa equipment at Surrey.  WNC

(b) BRF reported that Kui Zhao, who is Head of Nuclear Physics in Beijing and has recently spent time at Manchester, had approached the group. She was interested in discussing a joint agreement for exchanges.

16. Date of Next Meetings:

Next Meeting: Thursday, 13 May 1993 in Birmingham,

The date of the subsequent meeting was also decided. ^{‡}

W.N. Catford, Secretary. April 26, 1993 (released May 4, 1993)

‡Following Meeting: Thursday, 15 July 1993 in Oxford. (BRF, GT and SJAH will be in Canberra)